OpenStudy (anonymous):

Integrate the integral below:

OpenStudy (anonymous):

\[\int\limits\limits_{}^{}\frac{ xdx }{ (5-2x)^{\frac{ 3 }{2 }} }\] Looking at it, I think the only way to solve the integral above is by using substitution. So I let \[t= (5-2x)^{\frac{ 3 }{ 2 }}\] from this I got the value of \[x= \frac{ 5-t ^{\frac{ 2 }{ 3 }} }{ 2 }\] which is \[dx= -\frac{ dt }{ 3t ^{\frac{ 1 }{ 3 }} }\] ^I am not very sure if I got the derivative right Then I need to substitute the values to the original integral, which will be equal to \[\int\limits\limits_{}^{}\frac{ \frac{ 5dt-t ^{\frac{ 2 }{ 3 }}dt }{ 6t ^{\frac{ 1 }{ 3 }} } }{ t }\] Then after this I don't know what to do anymore, I'm not even sure if I am on the right track. Help?

OpenStudy (anonymous):

Try \[t = (5-2x)^{1/2}\]

OpenStudy (anonymous):

why did you use 1/2 instead of 3/2?

OpenStudy (anonymous):

Let me calculate in my rough notebook.

OpenStudy (anonymous):

ok, I'll wait :) But what is the reason on why you advised me to use 1/2?

sam (.sam.):

How about u=5-2x du=-2dx -du/2=dx \[\huge -\frac{1}{2} \int\limits \frac{5-u}{2 u^{\frac{3}{2}}} \, du\]

OpenStudy (anonymous):

no bro! I think I got it.

OpenStudy (anonymous):

\[\frac{dt}{dx} = \frac{-1}{(5-2x)^{1/2}} \]

sam (.sam.):

Split and integrate separately \[-\frac{1}{2} \int\limits \frac{5}{2 u^{\frac{3}{2}}}-\frac{1}{2 \sqrt{u}} \, du\]

OpenStudy (anonymous):

\[x = \frac{5-t^2}{2}\]

OpenStudy (anonymous):

integral becomes \[\int\limits_{}^{} \frac{(5-t^2) dt}{2t^2}\]

OpenStudy (anonymous):

Go ahead @taiga_aisaka

OpenStudy (anonymous):

Did you still used \[t= (5-2x)^{\frac{ 1 }{ 2 }}?\]

OpenStudy (anonymous):

Yes I did this using that substitution !

OpenStudy (anonymous):

Ok. Thank you very much.

OpenStudy (anonymous):

First tell that did you get the right answer ? :P lol

OpenStudy (anonymous):

Not yet, I'm currently solving it.